# The JWST-NIRCam View of Sagittarius C. III. The Extinction Curve

### v1

@@ 4.2 Consecutive filters method
The reference filter method always uses a fixed wavelength to estimate the extinction ratios. To check whether this approach misses any potential variation between consecutive wavelengths—and to validate the results obtained—we computed the extinction ratios between consecutive filters using the same approach described above. However, we did not derive the extinction ratios and because the differential extinction between these filter pairs is very small. This is not merely due to their proximity in wavelength, but rather because the extinction curve flattens significantly at longer wavelengths, leading to nearly constant extinction values (see Table 3). Consequently, the slope of the RC features in the F470N vs. F405N–F470N and F480M vs. F470N–F480M CMDs becomes excessively steep, with the RC feature exhibiting a very narrow spread in colour. In such cases, uncertainties in colour may dominate the observed slope, leading to unreliable estimates of slopes and thus extinction ratios. Moreover, the transmission curve of F480M overlaps with that of F470N.
It is also important to note that in CMDs such as F480M vs. F470N–F480M, which use only LW filters, the foreground population does not appear as a clearly separated component, unlike in other CMDs, but is instead blended with Sgr C stars. To address this, we identified and removed foreground stars using the F212N vs. F162M–F212N CMD. This removal had a negligible impact on our results, as only a small fraction of these stars belong to the RC features (see Appendix C).
The extinction curve derived using the consecutive filters method is shown in Fig. 4. To enable a direct comparison, we converted the extinction ratios obtained with this approach to the reference filter system. As before, the final extinction-ratio values and their associated uncertainties were computed as the mean between the values obtained for the bright and faint RC features. As shown in Fig. 4, both methods yield extinction curves that agree within the uncertainties.
Since the consecutive filters method does not allow us to derive extinction ratios for the two reddest filters, and given its agreement with the reference filter method, we adopt the extinction curve obtained with the latter for the remainder of the discussion.
@@ 5.1 Interpretation of the derived extinction curve
As shown in Fig. 4 and Tables 2 and 3, our extinction curve flattens and appears to keep slightly descending in the LW regime. However, the ratio appears higher than the trend suggested by neighboring filters in the LW regime. This deviation is visible as a bump in the interpolated extinction curve and is attributed to the strong CO ice absorption feature affecting the 4.7 m filter, as previously reported in Fig. 7(a) of Ginsburg et al. (2023) (see also Lutz, 1999; Moultaka et al., 2009).
@@
We observe a clear difference in the extinction indices derived for the SW filters (see Table 2) and those for the LW filters (see Table 3). In the SW range, we obtain , whereas in the LW range, the values decrease to . This indicates a clear wavelength dependence of the extinction index when comparing the short and long wavelength regimes. However, within the SW and LW regimes our measurements are consistent with single -values. This contrasts with previous studies such as Nogueras-Lara et al. (2020) and Hosek et al. (2018), which reported a noticeable wavelength dependence of the extinction index within the JH bands. However, our uncertainties are likely too large to detect such a variation.
5.2 Comparison with GALACTICNUCLEUS survey
@@
To further assess our results using an independent dataset and to investigate a possible wavelength dependence of the extinction index in the –m range, we repeated the analysis described in Sect. 4.1 using data from the GALACTICNUCLEUS survey (Nogueras-Lara et al., 2018, 2019b) that spatially overlaps with our JWST observations. The overlapping region is shown in Fig. 1. GALACTICNUCLEUS uses observations taken with the HAWK-I instrument at the VLT (Kissler-Patig et al., 2008). We selected these data because they are the most complete ones after JWST, thanks to their angular resolution, and because the GALACTICNUCLEUS JH filters overlap with our JWST filters—specifically, J with F115W, H with F162M, and with F212N. Due to the limited number of detections in the GALACTICNUCLEUS J band in this region (only 1700 stars), we restricted our analysis to the H and bands.
@@
We retrieved the GALACTICNUCLEUS data from the ESO Archive333https://archive.eso.org/scienceportal/home. Figure 10 shows the RC selection box and binning. We computed the extinction ratios for both RC components, along with their uncertainties, and derived the corresponding extinction indices using Eq. 4. By combining the statistical and systematic uncertainties in quadrature, and averaging the results from both RC components, we obtained a final extinction index of . This result is consistent with the value reported by Nogueras-Lara et al. (2020), , but it differs by about one sigma from the value obtained with our JWST data, . It is important to note that the region we analysed from GALACTICNUCLEUS contains RC stars, whereas Nogueras-Lara et al. (2020) used more than 165 000 RC stars in a much larger field, which may explain the smaller uncertainty of their index.
@@
To understand the discrepancy, we compared the transmission curves of the HAWK-I and JWST filters (see Fig. 5). The differences are considerable, particularly between the F212N and filters, with the HAWK-I bands being significantly broader. As a result, the effective wavelengths of the HAWK-I (m, m; Nogueras-Lara et al., 2018) and JWST (m, m) filters are different. If the extinction curve varies significantly within the 1–2.5 m range (as proposed by Hosek et al., 2018; Nogueras-Lara et al., 2019a, 2020), then even small shifts in effective wavelength can translate into measurable variations in the derived extinction index, which might justify the difference that we found.
@@
In addition, when using broad-band filters the measured extinction index and effective wavelength will depend on the type of stars observed. As discussed in Nogueras-Lara et al. (2020), this can introduce systematic biases.
Taken together, the combination of differences in filter profiles, the averaging effects of broad-band filters, and the presence of non-linear photometric effects provides a natural explanation for the discrepancy observed between our extinction index and that of the GALACTICNUCLEUS dataset. We cannot confirm the wavelength dependence of the extinction index in the 1–2.5 m range reported by Nogueras-Lara et al. (2020). However, considering the differences in the datasets, the targeted regions, and the challenges associated with accurately determining extinction, we cannot exclude this possibility either.
@@ 5.3 Comparison with previous extinction curves
Numerous infrared extinction curves have been reported in the literature, several of them based on GC data, obtained using different instruments and methodologies—each with its own wavelength coverage. As a result, a range of values for extinction ratios and indices has been published. In this section, we compare our derived extinction curve with previous curves obtained both from GC data and from observations targeting other regions within the Galaxy and, in one case, from an extragalactic environment.
@@ 5.3.1 Extinction curves derived with GC data
We contrasted our extinction curve (using the reference filter method) with several extinction curves from the literature which are representative for the GC. Figure 6 shows the comparison with the curves of Chiar & Tielens (2006), Nishiyama et al. (2009), and Fritz et al. (2011) (left), Schödel et al. (2010) and Hosek et al. (2019) (centre) and Alonso-García et al. (2017), Nogueras-Lara et al. (2018) and Nogueras-Lara et al. (2020) (right). To express all extinction curves relative to our reference filter, we interpolated them to estimate the extinction at the effective wavelength of our reference filter (F162M). For the interpolation, we used the CubicSpline function from the SciPy library. In addition, Table 4 shows the extinction ratios interpolated for our filter set based on the literature curves, alongside our results (we added quadratically the statistical and systematic uncertainties). It is important to note that Fig. 6 does not display these interpolated values, but rather the extinction ratios originally reported in the respective works, converted to our reference filter system. The curves appear continuous simply because we connected the reported data points for visualisation purposes.
@@
Overall, the trend observed in our extinction curve is consistent with that reported in previous studies: the extinction decreases with increasing wavelength until it reaches small values with local variations due to absorption by different molecules and ices, as clearly seen in Fig. 6. Based on Table 4, we find that the two curves covering the full wavelength range—those of Nishiyama et al. (2009) and Fritz et al. (2011)—are in good agreement with our results, except for in the case of Nishiyama et al. (2009).
@@
To reproduce the Nishiyama et al. (2009) extinction curve, we used the data provided in their Table 1. Their extinction curve was derived photometrically by combining Spitzer-IRAC observations of bulge fields with near-infrared SIRIUS-IRSF data of the GC, incorporating both space- and ground-based measurements. We reproduced the Fritz et al. (2011) curve using their publicly available interpolated infrared extinction curve. This curve was derived from hydrogen emission lines observed with ISO-SWS and SINFONI in the minispiral region.
@@
In both cases, the extinction curves correspond to regions that are spatially distinct from our Sgr C field: the Fritz et al. (2011) curve was derived in the central minispiral, while Nishiyama et al. (2009) analysed a broader area of the central bulge (, ). This supports previous evidence that the infrared extinction curve towards the inner Galaxy is not dependent on the line of sight (e.g. Nogueras-Lara et al., 2019b).
@@
Regarding the extinction index, Fritz et al. (2011) reported a mean value of over the wavelength range –m, while Nishiyama et al. (2009) found in the J, H, and bands. The latter is in better agreement with the extinction indices we derived for the SW filters. A plausible explanation for this difference lies in the methodologies used. Nishiyama et al. (2009) derived their extinction curve using a photometric method comparable to ours: they analysed RC stars in near-infrared colour-magnitude and colour-colour diagrams. In contrast, Fritz et al. (2011) used spectroscopic measurements. Thus, both our method and that of Nishiyama et al. (2009) rely on photometry of RC stars, sampling a similar stellar population and being subject to comparable biases—such as filter transmission effects.
@@
Our results are not consistent with those of Chiar & Tielens (2006), as shown in the left panel of Fig. 6 and in Table 4. The interpolated extinction ratios are systematically higher than ours. Chiar & Tielens (2006) combined a photometrically derived continuum extinction curve—based on the works of Lutz (1999) and Indebetouw et al. (2005)—with spectroscopic ISO-SWS data from GCS 3, an infrared-bright source in the Quintuplet cluster. However, robust conclusions are limited by the lack of reported uncertainties in their extinction values.
@@
Our extinction curve also shows good agreement with that of Schödel et al. (2010), which was derived from photometric observations of the central parsec of the Galaxy. This agreement is evident in both the central panel of Fig. 6 and Table 4. Their extinction indices, and , are consistent—within relatively large uncertainties, which likely stem from significant sources of error such as zero point uncertainties and the limited number of RC stars used (fewer than )—with our corresponding values: and . These particular filters were chosen due to their overlap. The first index was computed directly as described in Appendix B, while we calculated the second using Eq. 4, based on the derived extinction ratios and effective wavelengths. We obtained its associated uncertainty through standard quadratic error propagation.
@@
Regarding the extinction curve of Hosek et al. (2019) which spans the 0.7–3.54 m range, the authors combined HST observations of the GC (, ) and Westerlund 1 (, ) with VVV -band data for Westerlund 1, thereby integrating both space- and ground-based datasets and covering different lines of sight. The extinction ratios interpolated from their curve agree with our results within uncertainties, supporting the hypothesis that the infrared extinction curve towards the GC is not dependent on the line of sight.
@@
In contrast, our extinction curve is not compatible with those reported by Alonso-García et al. (2017), Nogueras-Lara et al. (2018), and Nogueras-Lara et al. (2020), as shown in the right panel of Fig. 6. Alonso-García et al. (2017) analysed several regions across the Galactic bulge using broad-band filters (Z, Y, J, H, ) from the VVV survey. Nogueras-Lara et al. (2018) and Nogueras-Lara et al. (2020) studied different areas within the GC using GALACTICNUCLEUS data. These studies employed the same methodology as ours to derive the extinction curve.
@@
To enable a more direct comparison, we interpolated our extinction curve to estimate the extinction at the J band using the effective wavelength computed in Nogueras-Lara et al. (2018) (this was also done for the central panel in Fig. 6). For the interpolation, we used the CubicSpline function. Visually, the value of aligns only with the result from Alonso-García et al. (2017). To further examine this, we computed for comparison with values reported in the literature, interpolating our curve to obtain . We derived . This is lower than the values reported by Alonso-García et al. (2017) (), Nogueras-Lara et al. (2018) (), and Nogueras-Lara et al. (2020) (). Although the value from Alonso-García et al. (2017) appears to be the closest, it still does not agree with our result within the uncertainties. The apparent agreement in Fig. 6 is misleading. Since the authors do not report effective wavelengths, we had to adopt them from the Spanish Virtual Observatory Filter Profile Service444http://svo2.cab.inta-csic.es/svo/theory/fps/ (Rodrigo et al., 2012; Rodrigo & Solano, 2020), introducing uncertainty in the x-axis placement of . Consequently, its true position could differ significantly from what is shown in the plot. The extinction ratio showing the largest disagreement is , which supports the idea of a wavelength-dependent extinction curve. The broad-band filters used in those studies are significantly wider than F212N, which is a narrow-band filter (see Fig. 5).
@@
Regarding the extinction indices, Alonso-García et al. (2017) found over their full wavelength range. Nogueras-Lara et al. (2018) reported for the range, and Nogueras-Lara et al. (2020) found two values: and . All of these are larger than the extinction indices we derived from the SW filters. In order to make decisive progress, we propose to map the extinction curve in the GC systematically with narrow band filters, from the space and from the ground and with space-based spectroscopy (NIRSpec and MIRI).

### v2

@@ 4.2 Consecutive filters method
The reference filter method always uses a fixed wavelength to estimate the extinction ratios. To check whether this approach misses any potential variation between consecutive wavelengths—and to validate the results obtained—we computed the extinction ratios between consecutive filters using the same approach described above. However, we did not derive the extinction ratios and because the differential extinction between these filter pairs is very small. This is not merely due to their proximity in wavelength, but rather because the extinction curve flattens significantly at longer wavelengths, leading to nearly constant extinction values (see Table 3). Consequently, the slope of the RC features in the F470N versus F405N–F470N and F480M versus F470N–F480M CMDs becomes excessively steep, with the RC feature exhibiting a very narrow spread in colour. In such cases, uncertainties in colour can dominate the observed slope, leading to unreliable estimates of slopes and thus extinction ratios. Moreover, the transmission curve of F480M overlaps with that of F470N.
It is also important to note that in CMDs such as F480M versus F470N–F480M, which use only LW filters, the foreground population does not appear as a clearly separated component, unlike in other CMDs, but is instead blended with Sgr C stars. To address this, we identified and removed foreground stars using the F212N versus F162M–F212N CMD. This removal had a negligible impact on our results, as only a small fraction of these stars belong to the RC features (see Appendix C).
The extinction curve derived using the consecutive filters method is shown in Fig. 4. To enable a direct comparison, we converted the extinction ratios obtained with this approach to the reference filter system. As before, the final extinction-ratio values and their associated uncertainties were computed as the mean between the values obtained for the bright and faint RC features. As shown in Fig. 4, the two methods yield extinction curves that agree within the uncertainties.
Since the consecutive filters method does not allow us to derive extinction ratios for the two reddest filters, and given its agreement with the reference filter method, we adopt the extinction curve obtained with the reference filter method for the remainder of the discussion.
@@ 5.1 Interpretation of the derived extinction curve
As shown in Fig. 4 and Tables 2 and 3, our extinction curve flattens and appears to keep slightly descending in the LW regime. However, the ratio appears higher than the trend suggested by neighbouring filters in the LW regime. This deviation is visible as a bump in the interpolated extinction curve and is attributed to the strong CO ice absorption feature affecting the 4.7 m filter, as previously reported in Fig. 7(a) of Ginsburg2023.
@@
We observe a clear difference in the extinction indices derived for the SW filters (see Table 2) and those for the LW filters (see Table 3). In the SW range, we obtain , whereas in the LW range, the values decrease to . This indicates a clear wavelength dependence of the extinction index when comparing the SW and LW regimes. However, within the SW and LW regimes our measurements are consistent with single -values. This contrasts with previous studies such as NoguerasLara2020 and Hosek2018, which reported a noticeable wavelength dependence of the extinction index within the JH bands. However, our uncertainties are likely too large to detect such a variation.
5.2 Comparison with the GALACTICNUCLEUS survey
@@
To further assess our results using an independent dataset and to investigate a possible wavelength dependence of the extinction index in the –m range, we repeated the analysis described in Sect. 4.1 using data from the GALACTICNUCLEUS survey (NoguerasLara2018; NoguerasLara2019) that spatially overlaps with our JWST observations. The overlapping region is shown in Fig. 1. GALACTICNUCLEUS uses observations taken with the HAWK-I instrument at the VLT (KisslerPatig2008). We selected these data because they are the most complete ones after JWST, thanks to their angular resolution, and because the GALACTICNUCLEUS JH filters overlap with our JWST filters—specifically, J with F115W, H with F162M, and with F212N. Due to the limited number of detections in the GALACTICNUCLEUS J band in this region (only 1700 stars), we restricted our analysis to the H and bands.
@@
We retrieved the GALACTICNUCLEUS data from the ESO Archive777https://archive.eso.org/scienceportal/home. Figure 10 shows the RC selection box and binning. We computed the extinction ratios for both RC components, along with their uncertainties, and derived the corresponding extinction indices using Eq. 3. By combining the statistical and systematic uncertainties in quadrature, and averaging the results from both RC components, we obtained a final extinction index of . This result is consistent with the value reported by NoguerasLara2020, , but it differs by about one sigma from the value obtained with our JWST data, . It is important to note that the region we analysed from GALACTICNUCLEUS contains RC stars, whereas NoguerasLara2020 used more than 165 000 RC stars in a much larger field, which may explain the smaller uncertainty of their index.
@@
To understand the discrepancy, we compared the transmission curves of the HAWK-I and JWST filters (see Fig. 5). The differences are considerable, particularly between the F212N and filters, with the HAWK-I bands being significantly broader. As a result, the effective wavelengths of the HAWK-I (m, m; NoguerasLara2018) and JWST (m, m) filters are different. If the extinction curve varies significantly within the 1–2.5 m range (as proposed by Hosek2018; NoguerasLara2019a; NoguerasLara2020), then even small shifts in effective wavelength can translate into measurable variations in the derived extinction index, which might justify the difference that we found.
@@
In addition, when using broad-band filters the measured extinction index and effective wavelength will depend on the type of stars observed. As discussed in NoguerasLara2020, this can introduce systematic biases.
Taken together, the combination of differences in filter profiles, the averaging effects of broad-band filters, and the presence of non-linear photometric effects provides a natural explanation for the discrepancy observed between our extinction index and that of the GALACTICNUCLEUS dataset. We cannot confirm the wavelength dependence of the extinction index in the 1–2.5 m range reported by NoguerasLara2020. However, considering the differences in the datasets, the targeted regions, and the challenges associated with accurately determining extinction, we cannot exclude this possibility either.
@@ 5.3 Comparison with previous extinction curves
Numerous infrared extinction curves have been reported in the literature, several of them based on GC data, obtained using different instruments and methodologies—each with its own wavelength coverage. As a result, a range of values for extinction ratios and indices has been published. In this section we compare our derived extinction curve with previous curves obtained both from GC data and from observations targeting other regions within the Galaxy and, in one case, from an extragalactic environment.
@@ 5.3.1 Extinction curves derived with GC data
We contrasted our extinction curve (using the reference filter method) with several extinction curves from the literature that are representative of the GC (see Fig. 6). To express all extinction curves relative to our reference filter, we interpolated them to estimate the extinction at the effective wavelength of our reference filter (F162M). For the interpolation, we used the CubicSpline function from the SciPy library. In addition, Table 4 shows the extinction ratios interpolated for our filter set based on the literature curves, alongside our results (we added quadratically the statistical and systematic uncertainties). It is important to note that Fig. 6 does not display these interpolated values, but rather the extinction ratios originally reported in the respective works, converted to our reference filter system. The curves appear continuous simply because we connected the reported data points for visualisation purposes.
@@
Overall, the trend observed in our extinction curve is consistent with that reported in previous studies: the extinction decreases with increasing wavelength until it reaches small values with local variations due to absorption by different molecules and ices, as clearly seen in Fig. 6. Based on Table 4, we find that the two curves covering the full wavelength range—those of Nishiyama2009 and Fritz2011—are in good agreement with our results, except for in the case of Nishiyama2009.
@@
To reproduce the Nishiyama2009 extinction curve, we used the data provided in their Table 1. Their extinction curve was derived photometrically by combining Spitzer Infrared Array Camera (IRAC; Fazio2004) observations of bulge fields with Simultaneous-Color InfraRed Imager for Unbiased Survey of the InfraRed Survey Facility (SIRIUS-IRSF; Nagashima1999) data of the GC, incorporating both space- and ground-based measurements. We reproduced the Fritz2011 curve using their publicly available interpolated infrared extinction curve. This curve was derived from hydrogen emission lines observed with ISO-SWS and SINFONI in the mini-spiral region.
@@
In both cases, the extinction curves correspond to regions that are spatially distinct from our Sgr C field: the Fritz2011 curve was derived in the central mini-spiral, while Nishiyama2009 analysed a broader area of the central bulge (, ). This supports previous evidence that the infrared extinction curve towards the inner Galaxy is not dependent on the line of sight (e.g. NoguerasLara2019).
@@
Regarding the extinction index, Fritz2011 reported a mean value of over the wavelength range –m, while Nishiyama2009 found in the J, H, and bands. The is in better agreement with the extinction indices we derived for the SW filters. A plausible explanation for this difference lies in the methodologies used. Nishiyama2009 derived their extinction curve using a photometric method comparable to ours: they analysed RC stars in near-infrared colour-magnitude and colour-colour diagrams. In contrast, Fritz2011 used spectroscopic measurements. Thus, both our method and that of Nishiyama2009 rely on photometry of RC stars, sampling a similar stellar population and being subject to comparable biases—such as filter transmission effects.
@@
Our results are not consistent with those of Chiar2006, as shown in the left panel of Fig. 6 and in Table 4. The interpolated extinction ratios are systematically higher than ours. Chiar2006 combined a photometrically derived continuum extinction curve—based on the works of Lutz1999 and Indebetouw2005—with spectroscopic ISO-SWS data from GCS 3, an infrared-bright source in the Quintuplet cluster. However, robust conclusions are limited by the lack of reported uncertainties in their extinction values.
@@
Our extinction curve also shows good agreement with that of Schoedel2010, which was derived from photometric observations of the central parsec of the Galaxy. This agreement is evident in both the central panel of Fig. 6 and Table 4. Their extinction indices, and , are consistent—within relatively large uncertainties, which likely stem from significant sources of error such as zero point uncertainties and the limited number of RC stars used (fewer than )—with our corresponding values: and . These particular filters were chosen due to their overlap. The first index was computed directly as described in Appendix B, while we calculated the second using Eq. 3, based on the derived extinction ratios and effective wavelengths. We obtained its associated uncertainty through standard quadratic error propagation.
@@
Regarding the extinction curve of Hosek2019, which spans the 0.7–3.54 m range, the authors combined Hubble Space Telescope observations of the GC (, ) and Westerlund 1 (, ) with VISTA Variables in the Via Lactea (VVV; Minniti2010) -band data for Westerlund 1, thereby integrating both space- and ground-based datasets and covering different lines of sight. The extinction ratios interpolated from their curve agree with our results within uncertainties, supporting the hypothesis that the infrared extinction curve towards the GC is not dependent on the line of sight.
@@
In contrast, our extinction curve is not compatible with those reported by AlonsoGarcia2017, NoguerasLara2018, and NoguerasLara2020, as shown in the right panel of Fig. 6. AlonsoGarcia2017 analysed several regions across the Galactic bulge using broad-band filters (Z, Y, J, H, and ) from the VVV survey. NoguerasLara2018 and NoguerasLara2020 studied different areas within the GC using GALACTICNUCLEUS data. These studies employed the same methodology as ours to derive the extinction curve.
@@
To enable a more direct comparison, we interpolated our extinction curve to estimate the extinction at the J band using the effective wavelength computed in NoguerasLara2018. For the interpolation, we used the CubicSpline function. Visually, the value of aligns only with the result from AlonsoGarcia2017. To further examine this, we computed for comparison with values reported in the literature, interpolating our curve to obtain . We derived . This is lower than the values reported by AlonsoGarcia2017, NoguerasLara2018, and NoguerasLara2020. Although the value from AlonsoGarcia2017 appears to be the closest, it still does not agree with our result within the uncertainties. The apparent agreement in Fig. 6 is misleading. Since the authors do not report effective wavelengths, we had to adopt them from the Spanish Virtual Observatory Filter Profile Service999http://svo2.cab.inta-csic.es/svo/theory/fps/ (Rodrigo2012; Rodrigo2020), introducing uncertainty in the x-axis placement of . Consequently, its true position could differ significantly from what is shown in the plot. The extinction ratio showing the largest disagreement is , which supports the idea of a wavelength-dependent extinction curve. The broad-band filters used in those studies are significantly wider than F212N, which is a narrow-band filter (see Fig. 5).
@@
Regarding the extinction indices, AlonsoGarcia2017 found over their full wavelength range. NoguerasLara2018 reported for the range, and NoguerasLara2020 found two values: and . All of these are larger than the extinction indices we derived from the SW filters. To make decisive progress, we propose mapping the extinction curve in the GC systematically with narrow-band filters, from the space and from the ground and with space-based spectroscopy (NIRSpec and MIRI).